2/36 what you rolled the first time has nothing to do with the next roll. It's an independent event.
What is the probability of pulling a queen out of a shuffled card deck?
There is a 1 (one) in 13 (thirteen) chance that you will pull out a queen from a deck of 52 (fifty-two) cards.
3/12*3/11 = 9/132, or 6.818%.
What is a simple example of an empirical probability?
Emily got 12 texts and 5 were from jasmine. The probability is 5/12.
When rolling two six sided number cubes the probability that on one roll you get at least a nine?
It is 5/18.
Does something that can never happen have a probability of 0?
Yes, a probability of 0 is something that cannot happen.
Why normal distribution used in solving probability problems?
Because of the Central Limit Theorem, the mean value of sets of observations is distributed Normally. Irrespective of the underlying distribution of a variable, the distribution of a sum (or mean) of such variables tends towards the Normal distribution. So, for sufficiently large samples, all distributions can be approximated by thye Normal distribution.
One of the consequences is that the distribution has been studied extensively and a lot is known about tests based on it.
It is 0.4
Why are important in random variables?
It might help if you specified why WHAT was important in random variables.
What is the definition of probability?
The likelihood that a particular event will occur is called probability
The ratio of the wanted outcomes divided by the possible outcomes gives a fraction or ratio usually expressed as a percentage.
The likelihood of an event occuring.
What is an example of probability?
If you have 10 marbles in a bag, 5 red 2 blue, and 3 purple, and you randomly draw one out the percentage of it being Red=5/10=1/2 Blue=2/10=1/5 Purple=3/10
Number of ways 5 objects can be arranged in a row?
5! is the symbol for Factorial 5 , ie 5 x 4 x 3 x 2 = 120 different ways
The amount of time that a certain type of battery functions is a random variable?
A six-sided die, in which each side is equally likely to appear, is repeatedly rolled until the total of all rolls exceeds 400. Approximate the probability that this will require more than 140 rolls.
How many outcomes can you get if you make seven number random phone numbers?
The answer depends on what you are measuring. The random variable could be:
Other variables will generate other outcomes.
True or false p lies in plane b the line containing p and q must lie in plane b?
False.
In order for the line PQ to lie in plane B, then both P and Q must lie in plane B.
The answer to the question, as stated, is two. You are drawing from the same regular deck. If the questioner meant same suit, the answer is 5. The first four could be one of each of the four suits but the fifth must belong to one of the suits you have already drawn.
If two numbers are complementary what is their probability?
On the basis that numbers are a continuous variable, the probability of any particular number (or pair) is 0.